OpenStudy (lilday12):

A rock dropped on the moon will increase its speed from 0 m/s to 8.15 m/s in about 5 seconds. What is the acceleration of the rock? A. 8.2 m/s2 toward the moon surface B. 9.8 m/s2 toward the moon surface C. 41 m/s2 toward the moon surface D. 1.6 m/s2 toward the moon surface

OpenStudy (shikamaru11):

well according to the 1st eq. of motion is Vf=Vi+at so we can deduce a=Vf-Vi/t so just put the values for Vi = 0 as the initial is zero and final is 8.15m/s so here is your answer =1.6 m/s